# IV. Characterizing atmosphere gray extinction with thermal infrared observations

### v1

@@ Abstract
Context. Next-generation ground-based surveys, like the Legacy Survey of Space and Time at the Vera C. Rubin Observatory, demand highly precise calibration that remains stable over years and uniformly accurate across the sky to better than one percent, even under non-photometric conditions. This motivates new developments in photometric calibration, encompassing those designed to improve Earth’s atmosphere transmission characterization.
Aims. The StarDICE experiment seeks to establish a metrology chain linking laboratory standards to astrophysical fluxes, targeting 1 mmag accuracy in the griz bands. Reaching this precision requires mitigating variable atmospheric effects, especially gray extinction from clouds, which remains a major challenge. To address this challenge, we present a novel method for correcting photometric data using simultaneous radiometric measurements from an infrared thermal camera.
Methods. The gray extinction correction model is fitted on an image-by-image basis as a function of thermal radiance excess and difference between synthetic and instrumental fluxes of calibration stars. This approach stands out by avoiding any assumptions about spatial structure and leveraging a forward model combining in situ environmental monitoring, radiative transfer simulations and use of Gaia Data Release 3 star catalogs.
Results. We demonstrate the method using data collected from a dedicated experimental remote observation system, built to conduct repeated observations of two fields over several nights in varying atmosphere conditions. We show that applying the correction model to each source in the test sample results in a reduction of residuals between corrected and reference magnitudes. It delivers extinction correction maps for individual images with a resolution of 2 arcmin and an accuracy of $\sim$0.01 mag. This leads to a significant improvement in accuracy per image, with the mean absolute error decreasing from 0.64 to 0.11 mag for the most gray extinguished exposures. The temporal variations in extinction for non-ideal photometric sequences can be reduced to 0.025 mag per source, which is promising for its application in StarDICE.
Conclusions. We highlight the potential of this method for ground-based telescopes, discuss its implications for future survey strategies, and suggest possible avenues for improvement.
Key Words.:
atmospheric effects – radiative transfer – instrumentation: detectors – methods: data analysis – techniques: photometric
@@ 1 Introduction
The second way consists in exploring methods to correct for cloud extinction by fitting residuals in source photometry. Wang et al. (2012) analyzed data from the Chinese Small Telescope ARray (CSTAR; see Yuan et al. 2008) by constructing a two-dimensional interpolation map of gray extinction using a modified inverse distance weighting algorithm (Shepard, 1968). They found that over 80% of images were taken under even photometric conditions (extinction unevenness << 0.01 mag), while about 8% showed significant extinction variations (>> 0.02 mag). The method was effective in reducing dispersion in non-photometric sequences but showed no improvement under good conditions, and its global impact was not assessed. Ivezić et al. (2007) calibrated photometric zero points for non-photometric SDSS (York et al., 2000) data using dense stellar fields, achieving 2% accuracy even through 1 mag thick clouds. They suggested that further improvements could be obtained by modeling cloud opacity across multiple calibration patches, projecting that LSST could reach 1% calibration accuracy through 1–3 mag clouds. Building on this, Burke et al. (2013) applied two-dimensional polynomial fits to MOSAIC II data and achieved $\sim$0.5% precision in clear conditions and $\sim$1% repeatability through 1.5 mag thick clouds. They also quantified the spatial structure of cloud absorption, noting that sub-percent precision is limited by the presence of thin cloud layers. Finally, Burke et al. (2017) introduced the Forward Global Calibration Method (FGCM) for the Dark Energy Survey (Dark Energy Survey Collaboration et al., 2016), obtaining 6–7 mmag residual calibration errors per exposure. Although gray extinction from clouds was not explicitly modeled due to mostly clear observing conditions, FGCM estimated gray extinction as residuals between predicted and observed magnitudes, averaged over each CCD. More recently, Garrappa, S. et al. (2025) introduced a method of absolute calibration that simultaneously treats instrumental and atmospheric effects on an image-by-image basis by fitting the system transmission. They proceed in a similar manner to Burke et al. (2013) by fitting zero-point distribution using two-dimensional polynomials on sub-divisions of each single image and obtain an accuracy of the zero-point between 3–5 mmag for each exposure when applied on data from the Large Array Survey Telescope (LAST, see Ofek et al. 2023). All these methods rely on high-density stellar fields within the instrument’s FOV and show diminished performance when stellar density is low or when gray extinction exceeds $\sim$1.5 mag (Burke et al., 2013).
@@
We investigate here another approach that has been proposed in the conclusions of Burke et al. (2013) and Hazenberg (2019), which consists in collecting independent measurements of atmospheric down-welling radiance using an infrared thermal camera. We present our method to correct optical photometry from cloud extinction using thermal infrared observations in Section 2. We then present the instrumental setup in Section 3, and detail the observation campaign and collected data in Section 4. Section 5 and 6 are respectively devoted to the presentation of the analysis and results. Finally, in Section 7, we summarize the main findings of this study, and we discuss current limitations and potential improvements.
@@ 5.1 Selection criteria and catalog of objects
All stars meeting these selection criteria for each field were randomly divided into two subsets following a train-test split scheme. The first subset, called the “training” set, contains 70% of the sources used to adjust the parameters of the gray extinction models. These are the stellar calibrators. The second subset, called the “test” set, contains the remaining 30% of the sources and is used to evaluate the accuracy of the model.
@@ 6 Results
For the majority of exposures, there is either no gray extinction or the data is too sparse to reliably fit the non-linear deviation that appears at high values of gray extinction $\Delta m_{s,i}$. Instead, only a linear relationship between gray extinction and radiance excess is observable for nearly half of them. As a result, fitting the full physical model often yields unstable behavior, with parameters exhibiting unrealistically large covariances. To address this issue for images acquired under low but non-zero levels of gray extinction, we instead fit a simple linear model (Eq. 14) – with a positively bounded slope – whenever the relative uncertainty of either parameter $A$ or $B$ in the physical model exceeds 50%. If the fitted slope hits the lower bound of zero (or the uncertainty straddles around zero), no correction is applied to the observed magnitudes, as this indicates an absence of gray extinction in the image. Overall, approximately 5% of the images are corrected using the physical model described in Equations 1 and 13 (group n°1), $\sim$48% are corrected by the linear model (group n°2) and the remaining $\sim$47% do not contain sufficient gray extinction to be corrected for (group n°3). To illustrate the behavior of these three groups, we selected three exposures representing the three group types described above. Diagnostic plots for these cases are shown in Figures 6, 7, and 8. Top-left panels show the map of radiance excesses with black dots indicating the stars positions and the white rectangle the footprint of the CMOS image. Top-right panels show the measured gray extinction values for each star of the training (black errorbars) and test (gray errorbars) subsets plotted against the radiance excesses $\Delta L_{s,i}$ with the best-fit curve in red. The lower left and right panels of each figure shows respectively the estimated and inferred gray extinction values from the fit for test sources as a function of CMOS detector coordinates. The numbers indicated in the upper-left corner are the calculated dispersion of the gray extinction values.
@@ 6.1 Correction accuracy
To quantify the accuracy of the correction models, we estimate the propagated error on corrected magnitudes using the jacobi [^18] https://github.com/hDembinski/jacobi package. The function jacobi.propagate computes $\bm{C^{\prime}}=\bm{J}\bm{C}\bm{J}^{T}$, where $\bm{C}$ is the covariance matrix of the fitted parameters, $\bm{J}$ is the Jacobi matrix of first derivatives of the model computed numerically and $\bm{C^{\prime}}$ is the output matrix of the model prediction containing uncertainties. The distributions of the propagated uncertainties of the linear (green curve) and physical models (red curve) corrections are shown in Fig. 9 and are clipped to a minimum and maximum boundary at 1 mmag and 0.15 mag respectively. Values outside of this range are clipped to the given interval edges. They both have a Gamma distribution shape with values ranging from 1 mmag to 15 mmag for the high-end tail. The mean values are 5 mmag and 29 mmag for the linear and the physical model respectively. This can be explained by the fact that the physical model is used to correct for large extinctions. When examining the percentage ratio $\sigma^{\text{model}}/\sigma^{\text{phot}}$ between models propagated error and the photometric noise for each source (Eq. 4), both curves peak at 5% and the mean values are 10% and 21% for the linear and physical models respectively. This emphasizes the dominance of photon noise in the total uncertainty budget and the accuracy of the linear model being superior to the physical model.
@@ Appendix A Observation campaign
Table 1: Fields observed during the observation campaign.191919(*) Only these target fields were selected for the analysis. The following quantities are listed: (1) target name or identified; (2) and (3) right ascension and declination coordinates at J2000 epoch; (4) and (5) number of photometric and radiometric exposures; (6) number of nights/sequences; (7) number of remaining stellar objects after quality cuts described in Sec. 5.1.
@@
Polaris 02h31m49s 8915’51” 173 12430 2 698
HD 183030 17h16m56s 8902’16” 834 63012 5 724
HD 180609 19h12m47s 6410’37” 700 52741 3 1235
GD153 12h57m02s 2201’53” 450 33832 1 330
BD+28 4211 21h51m11s 2851’50” 6996 479195 11 466
HD 210955 22h13m22s 3122’56” 6918 458225 12 476
Table 2: Summary of the observation sequences of BD+28 4211.202020(*) This night is used as the reference sequence to compute the reference magnitudes $m_{s}^{\text{TOA, ref}}$. The following quantities are listed: (1) date of the sequence; (2) and (3) UTC start and end time of the sequence; (4) empirical qualitative assessment of atmospheric conditions; (5) and (6) number of photometric and radiometric exposures.
@@
Table 3: Summary of the observation sequences of HD 210955.212121(*) This night is used as reference the sequence to compute the reference magnitudes $m_{s}^{\text{TOA, ref}}$. The columns are identical to Table 2.
@@ Appendix B Ancillary data
  • Ozone total column and ground surface albedo are retrieved from the ERA-5222222https://cds.climate.copernicus.eu/datasets/reanalysis-era5-single-levels?tab=overview (Hersbach et al. 2020) dataset through the cdsapi [^23] https://cds.climate.copernicus.eu/how-to-api package and by applying linear interpolation in space and time for each observation epoch.
  • Aerosol properties: aerosol optical depth and the Ångström exponent are sourced from the AERONET [^24] https://aeronet.gsfc.nasa.gov/ (Holben et al. 1998) instrument at OHP which only provides daytime measurements. These parameters values are estimated by linear interpolation to the nighttime observation epochs.
  • Atmosphere vertical profiles: temperature, humidity, and ozone profiles are derived from radiosonde data collected at OHP [^25] Publicly available for download from the Network for the Detection of Atmospheric Composition Change (NDACC) : https://ndacc-aeris.ipsl.fr. Seasonal averages are computed with 2.5-$\sigma$ clipping from data spanning 2008–2022, and these profiles are supplemented by the US1976 standard model (Atmosphere 1976) for altitudes above 35,000 m, where atmospheric radiance and extinction contributions are minimal.
Table 4: Input parameters for simulations of optical transmission and infrared radiance with libRadtran.262626(*) Required to compute the cloud-free atmosphere down-welling radiance. (**) Required to compute the chromatic component of optical transmission.
@@ C.1 Airmass
Accurate airmass calculations based on the positions of sources are performed using the model by Young (1994). It provides an approximate formula to calculate the airmass as a function of the true zenith angle, rather than the refracted apparent zenith angle. The error in airmass calculated with this method is less than 0.001 up to $X=6$, with a maximum of 0.0037 at the horizon (Young 1994).
@@ C.2 Molecular extinction
Molecular Rayleigh scattering, along with absorption by oxygen and trace gases, is accurately quantified using barometric pressure data (Hansen & Travis 1974; Stubbs et al. 2007).
Atmospheric pressure varies slowly on timescales of several hours ($\pm$ 1 hPa) and precise local measurements are easily done with an in situ weather station. By combining precise knowledge of both the airmass (Section C.1) and the barometric pressure ($\pm$ 0.2 hPa), it is possible to provide a determination of molecular extinction with a precision of the order of a milli-magnitude on the observed stellar fluxes (Stubbs et al. 2007).
Ozone absorption can be described by a single parameter that is the integrated vertical column height from satellite-based measurements. This height is commonly expressed in Dobson Units (1 DU = 0.01 mm) and exhibits seasonal changes with minor fluctuations over several days (Merienne et al. 1989). Consequently, linear interpolation of hourly total ozone column is sufficient.
@@ C.3 Water vapour absorption
Water vapour absorption can be accurately simulated when the precipitable water vapour (PWV) is known. However, PWV varies rapidly over time [^27] After analyzing the daily data collected over several months, we observed PWV fluctuations of up to 30% on sub-hour timescales at OHP., and since H2O line strengths are highly sensitive to it, water vapour absorption significantly impacts photometric measurements, particularly in the $i$, $z$, and $y$ bands (see Fig. 13). Local PWV concentrations can evolve in the same way as aerosols, with variations of up to 10% per hour. This requires an evolving photometric correction. Variations of the order of $\Delta PWV$ = 1 mm affect the $z$ and $y$ band photometry of StarDICE by 5 and 25 mmag respectively. Measurements of water vapour content are made at 30 sec intervals using a dedicated ancillary instrument presented in Section 3.3.

### v2

# StarDICE IV: correcting visible photometry from atmospheric gray extinction using thermal infrared observationsD

@@ Abstract
Next-generation ground-based surveys such as the Vera C. Rubin Observatory’s Legacy Survey of Space and Time require photometric calibration that is both long-term stable and spatially uniform at the sub-percent level, even during non-photometric conditions. Achieving this precision motivates new approaches to characterize atmospheric transmission, particularly to mitigate gray extinction from clouds. The StarDICE experiment aims to establish a metrology chain linking laboratory standards to astrophysical fluxes with 1 mmag accuracy in the griz bands, a goal for which controlling variable atmospheric effects is essential. We present a method that corrects photometric measurements using simultaneous radiometric information from an infrared thermal camera. The gray-extinction model is fit on an image-by-image basis using thermal radiance excess and the difference between synthetic and instrumental fluxes of calibration stars, without requiring assumptions about the spatial structure of extinction. The method relies on a forward model that incorporates environmental monitoring, radiative-transfer simulations, and Gaia DR3 stellar catalogs. Using data from a remote observing system that repeatedly monitored two fields under diverse atmospheric conditions, we show that the corrections reduce residuals between corrected and reference magnitudes and produce extinction maps with 2-arcmin resolution and $\sim$0.01 mag accuracy. Using this technique, we can recover data acquired under non-photometric conditions with a precision comparable to data obtained under photometric conditions. For the most affected exposures, the mean absolute error improves from 0.64 to 0.11 mag, and temporal extinction variations can be reduced to 0.025 mag per source. We discuss the implications of this technique for future surveys and outline directions for further refinement.
keywords:
Instrumentation – Atmosphere – Transmission – Calibration – Photometry – Telescopes
@@ 1 Introduction
The second way consists in exploring methods to correct for cloud extinction by fitting residuals in source photometry. Wang et al. (2012) analyzed data from the Chinese Small Telescope ARray (CSTAR; see Yuan et al. 2008) by constructing a two-dimensional interpolation map of gray extinction using a modified inverse distance weighting algorithm (Shepard, 1968). They found that over 80 per cent of images were taken under even photometric conditions (extinction unevenness << 0.01 mag), while about 8 per cent showed significant extinction variations (>> 0.02 mag). The method was effective in reducing dispersion in non-photometric sequences but showed no improvement under good conditions, and its global impact was not assessed. Ivezić et al. (2007) calibrated photometric zero points for non-photometric SDSS (York et al., 2000) data using dense stellar fields, achieving 2 per cent accuracy even through 1 mag thick clouds. They suggested that further improvements could be obtained by modeling cloud opacity across multiple calibration patches, projecting that LSST could reach 1 per cent calibration accuracy through 1–3 mag clouds. Building on this, Burke et al. (2013) applied two-dimensional polynomial fits to MOSAIC II data and achieved $\sim$0.5 per cent precision in clear conditions and $\sim$1 per cent repeatability through 1.5 mag thick clouds. They also quantified the spatial structure of cloud absorption, noting that sub-percent precision is limited by the presence of thin cloud layers. Finally, Burke et al. (2017) introduced the Forward Global Calibration Method (FGCM) for the Dark Energy Survey (Dark Energy Survey Collaboration et al., 2016), obtaining 6–7 mmag residual calibration errors per exposure. Although gray extinction from clouds was not explicitly modeled due to mostly clear observing conditions, FGCM estimated gray extinction as residuals between predicted and observed magnitudes, averaged over each CCD. More recently, Garrappa, S. et al. (2025) introduced a method of absolute calibration that simultaneously treats instrumental and atmospheric effects on an image-by-image basis by fitting the system transmission. They proceed in a similar manner to Burke et al. (2013) by fitting zero-point distribution using two-dimensional polynomials on sub-divisions of each single image and obtain an accuracy of the zero-point between 3–5 mmag for each exposure when applied on data from the Large Array Survey Telescope (LAST, see Ofek et al. 2023). All these methods rely on high-density stellar fields within the instrument’s FOV and show diminished performance when stellar density is low or when gray extinction exceeds $\sim$1.5 mag (Burke et al., 2013).
@@
We investigate here another approach that has been proposed in the conclusions of Burke et al. (2013) and Hazenberg (2019) and which consists in gathering independent measurements of atmospheric down-welling radiance using an infrared thermal camera. Our goal is to maintain sufficient precision on individual observations collected under non-photometric conditions while maximizing the StarDICE telescope uptime. We present our method to correct optical photometry from cloud extinction using thermal infrared observations in Section 2. We then present the instrumental setup in Section 3, and detail the observation campaign and collected data in Section 4. Section 5 and 6 are respectively devoted to the presentation of the analysis and results. Finally, in Section 7, we summarize the main findings of this study, and we discuss current limitations and potential improvements.
@@ 5.1 Selection criteria and catalog of objects
All stars meeting these selection criteria for each field were randomly divided into two subsets following a train-test split scheme. The first subset, called the “training” set, contains 70 per cent of the sources used to adjust the parameters of the gray extinction models. These are the stellar calibrators. The second subset, called the “test” set, contains the remaining 30 per cent of the sources and is used to evaluate the accuracy of the model.
@@ 6 Results
For the majority of exposures, there is either no gray extinction or the data is too sparse to reliably fit the non-linear deviation that appears at high values of gray extinction $\Delta m_{s,i}$. Instead, only a linear relationship between gray extinction and radiance excess is observable for nearly half of them. As a result, fitting the full physical model often yields unstable behavior, with parameters exhibiting unrealistically large covariances. To address this issue for images acquired under low but non-zero levels of gray extinction, we instead fit a simple linear model (Eq. 14) – with a positively bounded slope – whenever the relative uncertainty of either parameter $A$ or $B$ in the physical model exceeds 50 per cent. If the fitted slope hits the lower bound of zero (or the uncertainty straddles around zero), no correction is applied to the observed magnitudes, as this indicates an absence of gray extinction in the image. Overall, approximately 5 per cent of the images are corrected using the physical model described in Equations 1 and  13 (group n°1), $\sim$48 per cent are corrected by the linear model (group n°2) and the remaining $\sim$47 per cent do not contain sufficient gray extinction to be corrected for (group n°3). To illustrate the behavior of these three groups, we selected three exposures representing the three group types described above. Diagnostic plots for these cases are shown in Figures 67, and 8. Top-left panels show the map of radiance excesses with black dots indicating the stars positions and the white rectangle the footprint of the CMOS image. Top-right panels show the measured gray extinction values for each star of the training (black errorbars) and test (gray errorbars) subsets plotted against the radiance excesses $\Delta L_{s,i}$ with the best-fit curve in red. The lower left and right panels of each figure shows respectively the estimated and inferred gray extinction values from the fit for test sources as a function of CMOS detector coordinates. The numbers indicated in the upper-left corner are the calculated dispersion of the gray extinction values.
@@ 6.1 Correction accuracy
To quantify the accuracy of the correction models, we estimate the propagated error on corrected magnitudes using the jacobi [^18] https://github.com/hDembinski/jacobi package. The function jacobi.propagate computes $\bm{C^{\prime}}=\bm{J}\bm{C}\bm{J}^{T}$, where $\bm{C}$ is the covariance matrix of the fitted parameters, $\bm{J}$ is the Jacobi matrix of first derivatives of the model computed numerically and $\bm{C^{\prime}}$ is the output matrix of the model prediction containing uncertainties. The distributions of the propagated uncertainties of the linear (green curve) and physical models (red curve) corrections are shown in Fig. 9 and are clipped to a minimum and maximum boundary at 1 mmag and 0.15 mag respectively. Values outside of this range are clipped to the given interval edges. They both have a Gamma distribution shape with values ranging from 1 mmag to 15 mmag for the high-end tail. The mean values are 5 mmag and 29 mmag for the linear and the physical model respectively. This can be explained by the fact that the physical model is used to correct for large extinctions. When examining the percentage ratio $\sigma^{\text{model}}/\sigma^{\text{phot}}$ between models propagated error and the photometric noise for each source (Eq. 4), both curves peak at 5 per cent and the mean values are 10 per cent and 21 per cent for the linear and physical models respectively. This emphasizes the dominance of photon noise in the total uncertainty budget and the accuracy of the linear model being superior to the physical model.
@@ Acknowledgements
Data Availability
The data underlying this article will be shared on reasonable request to the corresponding author.
@@ Appendix A Observation campaign
Table 1: Fields observed during the observation campaign. (*) Only these target fields were selected for the analysis. The following quantities are listed: (1) target name or identified; (2) and (3) right ascension and declination coordinates at J2000 epoch; (4) and (5) number of photometric and radiometric exposures; (6) number of nights/sequences; (7) number of remaining stellar objects after quality cuts described in Sec. 5.1.
@@
Polaris 02h31m49s 8915’51" 173 12430 2 698
HD 183030 17h16m56s 8902’16" 834 63012 5 724
HD 180609 19h12m47s 6410’37" 700 52741 3 1235
GD153 12h57m02s 2201’53" 450 33832 1 330
BD+28 4211 21h51m11s 2851’50" 6996 479195 11 466
HD 210955 22h13m22s 3122’56" 6918 458225 12 476
Table 2: Summary of the observation sequences of BD+28 4211. (*) This night is used as the reference sequence to compute the reference magnitudes $m_{s}^{\text{TOA, ref}}$. The following quantities are listed: (1) date of the sequence; (2) and (3) UTC start and end time of the sequence; (4) empirical qualitative assessment of atmospheric conditions; (5) and (6) number of photometric and radiometric exposures.
@@
Table 3: Summary of the observation sequences of HD 210955. (*) This night is used as reference the sequence to compute the reference magnitudes $m_{s}^{\text{TOA, ref}}$. The columns are identical to Table 2.
@@ Appendix B Ancillary data
  • Ozone total column and ground surface albedo are retrieved from the ERA-5191919https://cds.climate.copernicus.eu/datasets/reanalysis-era5-single-levels?tab=overview (Hersbach et al., 2020) dataset through the cdsapi [^20] https://cds.climate.copernicus.eu/how-to-api package and by applying linear interpolation in space and time for each observation epoch.
  • Aerosol properties: aerosol optical depth and the Ångström exponent are sourced from the AERONET [^21] https://aeronet.gsfc.nasa.gov/ (Holben et al., 1998) instrument at OHP which only provides daytime measurements. These parameters values are estimated by linear interpolation to the nighttime observation epochs.
  • Atmosphere vertical profiles: temperature, humidity, and ozone profiles are derived from radiosonde data collected at OHP [^22] Publicly available for download from the Network for the Detection of Atmospheric Composition Change (NDACC) : https://ndacc-aeris.ipsl.fr. Seasonal averages are computed with 2.5-$\sigma$ clipping from data spanning 2008–2022, and these profiles are supplemented by the US1976 standard model (Atmosphere, 1976) for altitudes above 35,000 m, where atmospheric radiance and extinction contributions are minimal.
Table 4: Input parameters for simulations of optical transmission and infrared radiance with libRadTran. (*) Required to compute the cloud-free atmosphere down-welling radiance. (**) Required to compute the chromatic component of optical transmission.
@@ C.1 Airmass
Accurate airmass calculations based on the positions of sources are performed using the model by Young (1994). It provides an approximate formula to calculate the airmass as a function of the true zenith angle, rather than the refracted apparent zenith angle. The error in airmass calculated with this method is less than 0.001 up to $X=6$, with a maximum of 0.0037 at the horizon (Young, 1994).
@@ C.2 Molecular extinction
Molecular Rayleigh scattering, along with absorption by oxygen and trace gases, is accurately quantified using barometric pressure data (Hansen & Travis, 1974; Stubbs et al., 2007).
Atmospheric pressure varies slowly on timescales of several hours ($\pm$ 1 hPa) and precise local measurements are easily done with an in situ weather station. By combining precise knowledge of both the airmass (Section C.1) and the barometric pressure ($\pm$ 0.2 hPa), it is possible to provide a determination of molecular extinction with a precision of the order of a milli-magnitude on the observed stellar fluxes (Stubbs et al., 2007).
Ozone absorption can be described by a single parameter that is the integrated vertical column height from satellite-based measurements. This height is commonly expressed in Dobson Units (1 DU = 0.01 mm) and exhibits seasonal changes with minor fluctuations over several days (Merienne et al., 1989). Consequently, linear interpolation of hourly total ozone column is sufficient.
@@ C.3 Water vapour absorption
Water vapour absorption can be accurately simulated when the precipitable water vapour (PWV) is known. However, PWV varies rapidly over time [^23] After analyzing the daily data collected over several months, we observed PWV fluctuations of up to 30 per cent on sub-hour timescales at OHP., and since H2O line strengths are highly sensitive to it, water vapour absorption significantly impacts photometric measurements, particularly in the $i$, $z$, and $y$ bands (see Fig. 13). Local PWV concentrations can evolve in the same way as aerosols, with variations of up to 10 per cent per hour. This requires an evolving photometric correction. Variations of the order of $\Delta PWV$ = 1 mm affect the $z$ and $y$ band photometry of StarDICE by 5 and 25 mmag respectively. Measurements of water vapour content are made at 30 sec intervals using a dedicated ancillary instrument presented in Section 3.3.